# MuSc-V2: Zero-Shot Multimodal Industrial Anomaly Classification and Segmentation with Mutual Scoring of Unlabeled Samples

### v1

@@ II-C Zero-shot anomaly classification and segmentation
In the industrial vision field, zero-shot AC/AS has garnered more attention. However, most methods focus on 2D modal by image-text alignment of CLIP [25]. These CLIP-based approaches [15, 38, 39, 40, 16, 41] fine-tune image encoder or text encoder to bridge the domain gap between natural and industrial scenarios. Meanwhile, some zero-shot methods [42, 17, 22] focus on detecting anomalies by using the unlabeled samples itself. [42] explores the relationship between patches inside one unlabeled image, but only handles texture products. ACR [17] proposes a new adaptation strategy without human involvement, which trains the network by other products from the dataset. For the 3D zero-shot task, PointAD [21] renders the point cloud to multiple images with different view angles. In this way, 2D methods could be used to process 3D data. In this paper, we propose a mutual scoring mechanism for both 2D/3D modals, which only uses the unlabeled samples without additional fine-tuning.
@@ II-E Manifold learning
In high-dimensional manifolds, Euclidean properties are only preserved in local space, making direct distance calculations unreliable. Some manifold learning methods [48, 49, 50, 51] construct an embedding space where distances align with the underlying manifold structure. Inspired by the above principles, we develop the RsCon module to refine pixel-level anomaly classification in manifolds.
@@ III Method
    Curvature Calculation: To correct point groups with discontinuous surfaces, we compute the curvature $\mathcal{C}$ at each group’s center point $p_{\textbf{c}}$ [54]. As shown in Fig.3 (a), we observe that when points from different surface (Surface-1) are incorrectly grouped, the curvature at $p_{\textbf{c}}$ increases significantly. Therefore, we perform the following steps for re-grouping if $\mathcal{C}$ exceeds a predefined threshold $\mathcal{C}_{thr}$.
@@
In the 2D case, we reshape the vectorized patch tokens $F_{\textbf{I}}^{i,s}$ into $\sqrt{M_{\textbf{I}}}\times\sqrt{M_{\textbf{I}}}\times C_{\textbf{I}}$ grid to restore the spatial position information for the convenience of neighbor searching. We extract $r\times r$ neighborhood features for a patch $m$ as $F_{\textbf{I}}^{i,s}(\mathcal{N}_{r}^{m})\in\mathbb{R}^{|\mathcal{N}_{r}^{m}|\times C_{\textbf{I}}}$, where $\mathcal{N}_{r}^{m}$ is the index of neighborhood patches and $|\mathcal{N}_{r}^{m}|$ is the number. Notably, a larger aggregation degree $r$ corresponds to a broader neighborhood, allowing the capture of larger anomaly regions. In the 3D case, due to the irregularity of point clouds, for each 3D patch, we identify its $r$ nearest neighborhood by computing the Euclidean distances between the current patch center and all other patches in the 3D space. The corresponding features of these neighboring patches are then defined as $F_{\textbf{P}}^{i,s}(\mathcal{N}_{r}^{m})\in\mathbb{R}^{|\mathcal{N}_{r}^{m}|\times C_{\textbf{P}}}$.
@@ III-C Multimodal Mutual Scoring
Our analysis above reveals that most unlabeled images in $\mathcal{D}$ assign lower scores to normal patches and higher scores to abnormal ones. Therefore, simple average operation on $A_{\textbf{I}}^{i,s,m}$ could effectively differentiate between normal and abnormal patches, shown in Fig. 5 (e). However, minor overlaps in (e) occur when some normal patches exhibit appearance variations across images, resulting in elevated scores. To mitigate this, we apply the Interval Average (IA) operation on the lowest $X\%$ of scores in $A_{\textbf{I}}^{i,s,m}$, suppressing outlier influences through:
@@ IV-A1 Datasets
We conduct experiments on two multimodal industrial datasets (MVTec 3D-AD [57] and Eyecandies [63]) and two 2D-only datasets (MVTec AD [64] and VisA [65]). MVTec 3D-AD consists of 10 product categories with 41 types of anomalies, where 3D points are stored as XYZ maps matching the RGB image resolution. Eyecandies provides synthetic data for 10 categories of candies, cookies and sweets, including depth maps and camera parameters for 3D reconstruction. For 2D-only datasets, MVTec AD contains 15 object/texture categories and VisA has 12 objects across 3 domains. All datasets provide normal and abnormal samples in the unlabeled test set. Notably, our method could operate effectively without requiring the full dataset. To validate its performance on the smaller dataset, we partition the original dataset into $g$ subsets and conduct anomaly detection independently within each subset.
@@ IV-A4 Competing methods
For 2D modal, we compare with some state-of-the-art zero-shot approaches, e.g. APRIL-GAN [16], AnomalyCLIP [52], ACR [17], AdaCLIP [38], VCP-CLIP [39] and FAPrompt [58]. For the CLIP-based methods, we load official checkpoints for inference. In multimodal scenarios, we evaluate against PointCLIPv2 [59], ULIP [60], ULIPv2 [61], PointAD [21], 3DzAL [62] and CMAD [66]. Since 3DzAL and CMAD have not been open-sourced, we only report the results from their official papers, with unavailable results denoted by ‘-’. In addition, we also compare few-shot methods, such as M3DM [18], CFM [47] and BTF [19].
TABLE II: Quantitative comparisons on MVTec AD and VisA datasets. Bold indicates the best performance. All metrics are in $\%$.
Dataset Method F1-max-cls AP-cls F1-max-seg AP-seg
MVTec AD AnomalyCLIP [52] 92.8 96.2 39.1 34.5
[64] AdaCLIP [38] 90.6 95.7 43.4 41.6
VCP-CLIP [39] 91.7 95.5 49.4 49.1
RareCLIP [67] 92.9 96.6 47.5 46.1
MuSc [22] 97.5 99.1 62.6 62.7
MuSc-V2 (Ours) 97.6 99.1 63.1 63.5
VisA AnomalyCLIP [52] 80.7 85.4 28.3 21.3
[65] AdaCLIP [38] 83.1 87.6 37.7 31.1
VCP-CLIP [39] 81.5 87.3 34.7 30.1
RareCLIP [67] 83.1 89.0 33.5 27.0
MuSc [22] 89.5 93.5 48.8 45.1
MuSc-V2 (Ours) 88.5 92.8 50.4 46.1
@@ IV-B Quantitative results
In Table I, we compare our MuSc-V2 with state-of-the-art zero-shot and few-shot methods on MVTec 3D-AD and Eyecandies datasets. We report the anomaly classification and segmentation results across 2D, 3D and multimodal settings. MuSc-V2 achieves superior performance in most metrics for all modals. Notably, it outperforms the second-best zero-shot method PointAD[21] by 23.7% and 19.3% AP for anomaly segmentation on these datasets. For anomaly classification, MuSc-V2 achieves $1.2\%$ and 6.2% AUROC gains on both datasets. When partitioning the original dataset into $g\in\{2,3\}$ subsets, both AC and AS metrics show minimal degradation ($\downarrow$): at most 1.0% on MVTec 3D-AD and 1.2% on Eyecandies. The slightly larger impact on Eyecandies stems from its limited sample size (50 samples per product). We report these performances as mean±std, with maximum standard deviations of 0.3 on MVTec 3D-AD and 0.7 on Eyecandies. These small variations confirm the insensitivity to different partitioning schemes and adaptability to diverse dataset compositions. When evaluated against 2D-focused zero-shot methods [52, 38, 39] on MVTec AD and VisA datasets (Table II), MuSc-V2 demonstrates significant improvements, even surpassing its previous version MuSc [22] across most metrics. However, some background regions on the VisA dataset contain subtle foreign impurities. Although these are not true anomalies, they are captured by the SNAMD module due to its sensitivity to subtle anomalies. This leads to over-detection, resulting in a slight performance drop on the VisA dataset.
@@ IV-D1 Effectiveness of the IPG strategy
TABLE IV: The ablation of SNAMD module. We report the anomaly classification and segmentation results. All metrics are in $\%$.
@@
IV-D2 Discussion of the SNAMD module
In Table IV, we conduct ablation experiments on two main technologies SWPooling and multiple degrees in our SNAMD module. The results demonstrate that three aggregation degrees $r\in\{1,3,5\}$ outperform using only one aggregation degree. It brings $0.8\%$ F1-max-cls and $1.6\%$ F1-max-seg improvements on the MVTec 3D-AD dataset, with consistent gains on Eyecandies. Removing SWPooling causes significant performance drops, particularly for the Eyecandies dataset with more small anomalies, where it reduce false negatives and brings $2.6\%$ AP-cls and $5.8\%$ AP-seg gains. When integrated into Local Neighborhood Aggregation (LNA) [55] and Local Neighborhood Aggregation with Multiple Degrees (LNAMD) [22], SWPooling also enhances their performance.
TABLE V: The ablation of four important modules in our multimodal MSM. We report the AC and AS results. All metrics are in $\%$.
Dataset Setting F1-max-cls AP-cls F1-max-seg AP-seg
MVTec 3D-AD w/o IA 92.4 96.1 52.5 52.5
w/o CAE 91.7 96.0 52.5 52.1
w/o $\lambda$ 92.4 96.7 54.4 54.3
Ours 93.0 96.8 54.6 54.7
Eyecandies w/o IA 80.7 83.6 41.6 38.4
w/o CAE 81.9 85.4 42.4 39.3
w/o $\lambda$ 82.2 85.9 43.7 40.8
Ours 82.9 86.1 44.7 41.8
IV-D3 Discussion of the Multimodal Mutual Scoring
Ablation studies in Table V evaluate four key components: Interval Average (IA), Cross-modal Anomaly Enhancement (CAE), and confidence weight $\lambda$. Without IA, normal regions with appearance variations receive higher scores from dissimilar patches, especially in Eyecandies where products have diverse sub-types. Our IA operation mitigates this issue, improving AP-cls by $2.5\%$ and AP-seg by $3.4\%$. The CAE module reduces false negatives brought by single-modal invisible anomalies, boosting MVTec 3D-AD performance by $1.3\%$ F1-max-cls and $2.1\%$ F1-max-seg. The confidence weight $\lambda$ in our CAE further suppresses cross-modal false positives, enhancing both classification and segmentation.
TABLE VI: The ablation of the RsCon module across four traditional datasets. We report AC and AS results. All metrics are in $\%$.
@@ IV-D5 Influence of the normal sample number
To investigate the robustness of the normal sample number in the test set, we randomly reduce normal samples to $\frac{1}{h}$ of the original set (Fig. 10). With $h=0$ as the limiting case (no normals), AS metrics show minimal degradation across both datasets: AUROC decreases by less than $0.23\%$ and PRO varies by less than $0.6\%$. The maximum observed drops in F1-max and AP are $2.94\%$, indicating only minor false-positive increases. These results demonstrate the insensitivity of our MuSc-V2 to normal sample counts, especially in real industrial scenes where normal samples typically dominate.
TABLE VII: Per sample inference time of our MuSc-V2 and other zero-shot methods. We divide the MVTec 3D-AD dataset into $g$ subsets.

### v2

@@ II-C Zero-shot anomaly classification and segmentation
In the industrial vision field, zero-shot AC/AS has garnered more attention. However, most methods focus on the 2D modal by image-text alignment of CLIP model [44]. These CLIP-based approaches [24, 6, 43, 33, 12, 19] fine-tune image encoder or text encoder to bridge the domain gap between the natural and industrial scenarios. Meanwhile, some zero-shot methods [1, 28, 32] focus on detecting anomalies by using the unlabeled samples itself. [1] explores the relationship between patches inside one unlabeled image, but only handles texture products. ACR [28] proposes a new adaptation strategy without human involvement, which trains the network by other products from the dataset. For the 3D zero-shot task, PointAD [66] renders the point cloud to multiple images with different view angles. In this way, 2D methods could be used to process 3D data.
Unlike general zero‑shot learning [26, 63] (non‑industrial one), where target domain data often contain diverse categories with large intra‑sample variations, industrial zero‑shot learning benefits from the inherent consistency of products manufactured on the same production line. This consistency means that unlabeled industrial samples share strong local similarities with each other. However, existing industrial zero‑shot methods typically overlook the rich relational information among unlabeled samples. Inspired by this observation, we propose a mutual scoring mechanism for both 2D/3D modals, which explicitly exploits the inter‑sample relationships within industrial unlabeled datasets without additional fine-tuning.
@@ II-E Manifold learning
In high‑dimensional feature spaces, the data often lie on a low‑dimensional manifold. Euclidean distances computed directly in the original space can be unreliable because they do not reflect the underlying manifold geometry, only local neighborhoods preserve Euclidean properties. To address this, some manifold learning methods [64, 34, 47, 61] construct an embedding space where distances approximate geodesic distances along the manifold. Inspired by the above principles, we develop the RsCon module to refine pixel-level anomaly classification by leveraging the local manifold structure. Each sample’s anomaly score is refined jointly with scores of its most similar neighbors on the manifold, ensuring that score calibration respects the intrinsic data geometry.
@@ III Method
    Curvature Calculation: To correct point groups with discontinuous surfaces, we compute the curvature $\mathcal{C}$ at each group’s center point $p_{\textbf{c}}$ [16]. Specifically, we define the $K_{\textbf{P}}$ points nearest to $p_{\textbf{c}}$ as $\{p_{1},...,p_{K_{\textbf{P}}}\}$. Then we normalize the coordinates of $p_{i}$ as follows,
    $\hat{p}_{i}=\frac{\overline{p}_{i}}{\text{max}(\|\overline{p}_{1}\|_{2},...,\|\overline{p}_{K_{\textbf{P}}}\|_{2})}$ (1)
    where $\overline{p}_{i}=p_{i}-p_{\textbf{c}}$. Furthermore, we calculate the eigenvalues $\{\lambda_{1},\lambda_{2},$...\} of the covariance matrix of the points $\{\hat{p}_{1},...,\hat{p}_{K_{\textbf{P}}}\}$. The curvature $\mathcal{C}$ is calculated by,
    $\mathcal{C}=\frac{\text{min}(\lambda_{1},\lambda_{2},...)}{\text{sum}(\lambda_{1},\lambda_{2},...)+\epsilon}$ (2)
    where $\epsilon$ is a small constant to prevent zero denominator. As shown in Fig.3 (a), we observe that when points from different surface (Surface-1) are incorrectly grouped, the curvature at $p_{\textbf{c}}$ increases significantly. Therefore, we perform the following steps for re-grouping if $\mathcal{C}$ exceeds a predefined threshold $\mathcal{C}_{thr}$.
@@
In the 2D case, we reshape the vectorized patch tokens $F_{\textbf{I}}^{i,s}$ into $\sqrt{M_{\textbf{I}}}\times\sqrt{M_{\textbf{I}}}\times C_{\textbf{I}}$ grid to restore the spatial position information for the convenience of neighbor searching. We extract $r\times r$ neighborhood features for a patch $m$ as $F_{\textbf{I}}^{i,s}(\mathcal{N}_{r}^{m})\in\mathbb{R}^{|\mathcal{N}_{r}^{m}|\times C_{\textbf{I}}}$, where $\mathcal{N}_{r}^{m}$ is the index list of $r\times r$ neighborhood patches of patch $m$, and $|\mathcal{N}_{r}^{m}|$ denotes its number. Notably, a larger aggregation degree $r$ corresponds to a broader neighborhood, allowing the capture of larger anomaly regions. In the 3D case, due to the irregularity of point clouds, for each 3D patch, we identify its $r$ nearest neighborhood by computing the Euclidean distances between the current patch center and all other patches in the 3D space. The corresponding features of these neighboring patches are then defined as $F_{\textbf{P}}^{i,s}(\mathcal{N}_{r}^{m})\in\mathbb{R}^{|\mathcal{N}_{r}^{m}|\times C_{\textbf{P}}}$.
@@ III-C Multimodal Mutual Scoring
Our analysis above reveals that most unlabeled images in $\mathcal{D}$ assign lower scores to normal patches and higher scores to abnormal ones. Therefore, simple average operation on all elements in the score list $A_{\textbf{I}}^{i,s,m}$ could effectively differentiate between normal and abnormal patches, shown in Fig. 5 (e). However, minor overlaps in (e) occur when some normal patches exhibit appearance variations across images, resulting in elevated scores. To mitigate this, we apply the Interval Average (IA) operation on the lowest $X\%$ of scores in $A_{\textbf{I}}^{i,s,m}$, suppressing outlier influences through:
@@ IV-A1 Datasets
We conduct experiments on two multimodal industrial datasets (MVTec 3D-AD [3] and Eyecandies [4]) and four 2D-only datasets (MVTec AD [2], VisA [71], MPDD [25], and BTAD [39]). MVTec 3D-AD consists of 10 product categories with 41 types of anomalies, where 3D points are stored as XYZ maps matching the RGB image resolution. Eyecandies provides synthetic data for 10 categories of candies, cookies and sweets, including depth maps and camera parameters for 3D reconstruction. For 2D-only datasets, MVTec AD contains 15 object/texture categories, VisA has 12 objects across 3 domains, the MPDD and BTAD datasets each contain 6 and 3 products respectively. All datasets provide normal and abnormal samples in the unlabeled test set. Notably, our method could operate effectively without requiring the full dataset. To verify its stability and robustness of on smaller datasets, we randomly partition the whole unlabeled dataset into $g$ subsets, each containing an equal number of samples, thereby simulating the behavior on small‑scale datasets. During evaluation, we collect the anomaly maps from all subsets and compute the metrics jointly to ensure a fair comparison.
@@ IV-A4 Competing methods
For 2D modal, we compare with some state-of-the-art zero-shot approaches, e.g. APRIL-GAN [12], AnomalyCLIP [65], ACR [28], AdaCLIP [6], VCP-CLIP [43] and RareCLIP [21]. For the CLIP-based methods, we load official checkpoints for inference. In multimodal scenarios, we evaluate against PointCLIPv2 [70], ULIP [55], ULIPv2 [56], PointAD [66], 3DzAL [50] and CMAD [38]. Since 3DzAL and CMAD have not been open-sourced, we only report the results from their official papers, with unavailable results denoted by ‘-’. In addition, we also compare few-shot methods, such as M3DM [51], CFM [14] and BTF [22].
TABLE II: Quantitative comparisons on four 2D-only datasets: MVTec AD, VisA, MPDD, and BTAD. Bold indicates the best performance.
Method Anomaly Classification Anomaly Segmentation Method Anomaly Classification Anomaly Segmentation
AUROC-cls F1-max-cls AP-cls AUROC-seg F1-max-seg AP-seg AUROC-cls F1-max-cls AP-cls AUROC-seg F1-max-seg AP-seg
MVTec AD [2] MPDD [25]
AnomalyCLIP [65] 91.5 92.8 96.2 91.1 39.1 34.5 AnomalyCLIP 77.0 79.9 82.0 96.5 34.2 28.9
AdaCLIP [6] 89.2 90.6 95.7 88.7 43.4 41.6 AdaCLIP 76.0 82.5 78.9 96.1 34.9 31.9
VCP-CLIP [43] 92.1 91.7 95.5 92.0 49.4 49.1 VCP-CLIP 69.3 78.1 72.9 96.8 31.2 27.8
RareCLIP [21] 91.5 92.9 96.6 91.5 47.5 46.1 RareCLIP 78.0 82.7 83.3 95.8 34.8 31.5
MuSc [32] 97.8 97.5 99.1 97.3 62.6 62.7 MuSc 78.2 83.8 79.3 97.5 36.3 33.0
MuSc-V2 (Ours) 97.7 97.6 99.1 97.3 63.1 63.5 MuSc-V2 (Ours) 79.0 83.6 78.7 97.5 37.7 32.6
VisA [71] BTAD [39]
AnomalyCLIP 82.1 80.7 85.4 95.5 28.3 21.3 AnomalyCLIP 88.3 83.8 87.3 94.2 49.7 45.5
AdaCLIP 85.8 83.1 87.6 95.5 37.7 31.1 AdaCLIP 88.6 88.2 93.3 92.1 51.7 47.8
VCP-CLIP 83.8 81.5 87.3 95.7 34.7 30.1 VCP-CLIP 89.3 88.8 92.6 93.8 48.5 45.1
RareCLIP 86.1 83.1 89.0 95.7 33.5 27.0 RareCLIP 91.6 89.1 93.4 91.7 54.1 50.1
MuSc 92.8 89.5 93.5 98.8 48.8 45.1 MuSc 94.8 94.6 97.5 97.3 57.5 54.1
MuSc-V2 (Ours) 91.9 88.5 92.8 98.8 50.4 46.1 MuSc-V2 (Ours) 96.5 96.1 99.1 97.5 59.0 57.3
@@ IV-B Quantitative results
In Table I, we compare our MuSc-V2 with state-of-the-art zero-shot and few-shot methods on MVTec 3D-AD and Eyecandies datasets. We report the anomaly classification and segmentation results across 2D, 3D and multimodal settings. MuSc-V2 achieves superior performance in most metrics for all modals. Notably, it outperforms the second-best zero-shot method PointAD[66] by 23.7% and 19.3% AP for anomaly segmentation on these datasets. For anomaly classification, MuSc-V2 achieves $1.2\%$ and 6.2% AUROC gains on both datasets. When partitioning the original dataset into $g\in\{2,3\}$ subsets, both AC and AS metrics show minimal degradation ($\downarrow$): at most 1.0% on MVTec 3D-AD and 1.2% on Eyecandies. The slightly larger impact on Eyecandies stems from its limited sample size (50 samples per product). We report these performances as mean±std, with maximum standard deviations of 0.3 on MVTec 3D-AD and 0.7 on Eyecandies. These small variations confirm the insensitivity to different partitioning schemes and adaptability to diverse dataset compositions. When evaluated against 2D-only zero-shot methods [65, 6, 43, 21] on four 2D-only datasets (Table II), MuSc-V2 demonstrates significant improvements, even surpassing its previous version MuSc [32] across most metrics. However, some background regions on the VisA dataset contain subtle noise. Although these are not true anomalies, they are captured by the SNAMD module due to its sensitivity to subtle anomalies. This affects the AC metric, while the AS metric remains nearly unchanged. The reason is that the false positives occur only in very small regions, resulting in a few pixels with high anomaly scores. This causes image‑level over‑detection but has only a negligible impact on pixel‑level evaluation. In addition, products in the MPDD dataset exhibit more pose variations, making it harder to find similar normal regions across images with different poses. While some existing methods achieve a slightly higher AP-cls, our method still delivers competitive or superior results on all other metrics.
@@ IV-D1 Effectiveness of the IPG strategy
TABLE IV: The ablation of four important modules in our multimodal MSM. We report the AC and AS results. All metrics are in $\%$.
Dataset Setting F1-max-cls AP-cls F1-max-seg AP-seg
MVTec 3D-AD w/o IA 92.4 96.1 52.5 52.5
w/o CAE 91.7 96.0 52.5 52.1
w/o $\lambda$ 92.4 96.7 54.4 54.3
Ours 93.0 96.8 54.6 54.7
Eyecandies w/o IA 80.7 83.6 41.6 38.4
w/o CAE 81.9 85.4 42.4 39.3
w/o $\lambda$ 82.2 85.9 43.7 40.8
Ours 82.9 86.1 44.7 41.8
IV-D2 Discussion of the Multimodal Mutual Scoring
Ablation studies in Table IV evaluate four key components: Interval Average (IA), Cross-modal Anomaly Enhancement (CAE), and confidence weight $\lambda$. Without IA, normal regions with appearance variations receive higher scores from dissimilar patches, especially in Eyecandies where products have diverse sub-types. Our IA operation mitigates this issue, improving AP-cls by $2.5\%$ and AP-seg by $3.4\%$. The CAE module reduces false negatives brought by single-modal invisible anomalies, boosting MVTec 3D-AD performance by $1.3\%$ F1-max-cls and $2.1\%$ F1-max-seg. The confidence weight $\lambda$ in our CAE further suppresses cross-modal false positives, enhancing both classification and segmentation.
TABLE V: The ablation of SNAMD module. We report the anomaly classification and segmentation results. All metrics are in $\%$.
@@ IV-D2 Discussion of the Multimodal Mutual Scoring
IV-D3 Discussion of the SNAMD module
In Table V, we conduct ablation experiments on two main technologies SWPooling and multiple degrees in our SNAMD module. The results demonstrate that three aggregation degrees $r\in\{1,3,5\}$ outperform using only one aggregation degree. It brings $0.8\%$ F1-max-cls and $1.6\%$ F1-max-seg improvements on the MVTec 3D-AD dataset, with consistent gains on Eyecandies. The $\exp$ operation amplifies the similarity weights of the similar neighborhood patches while suppressing those that are dissimilar. This effectively reduces the interference of irrelevant backgrounds, leading to consistent performance improvements on both datasets. Removing SWPooling causes significant performance drops, particularly for the Eyecandies dataset with more small anomalies, where it reduce false negatives and brings $2.6\%$ AP-cls and $5.8\%$ AP-seg gains. When integrated into Local Neighborhood Aggregation (LNA) [45] and Local Neighborhood Aggregation with Multiple Degrees (LNAMD) [32], SWPooling also enhances their performance.
IV-D4 Effectiveness of our RsCon module
[…]
TABLE VI: The ablation of the RsCon module across four traditional datasets. We report the AC results. All metrics are in $\%$.
@@ IV-D5 Influence of the normal sample number
To investigate the robustness of the normal sample number in the test set, we randomly reduce normal samples to $\frac{1}{h}$ of the original set (Fig. 10). The horizontal axis represents the value of $h$. The limiting case $h=0$ indicates no normal samples in the dataset. Under this extreme condition, AS metrics show minimal degradation across both datasets: AUROC decreases by less than $0.23\%$ and PRO varies by less than $0.6\%$. The maximum observed drops in F1-max and AP are $2.94\%$, indicating only minor false-positive increases. These results demonstrate the insensitivity of our MuSc-V2 to normal sample counts and the anomaly percentage, especially in real industrial scenes where normal samples typically dominate.
IV-D6 Effect of the dataset size
In our mutual scoring mechanism, we use $\mathcal{D}\backslash O_{i}\}$ to assign scores to sample $O_{i}$. To explore the sensitivity to dataset size, we divide the unlabeled samples into $g\in\{1,2,3\}$ subsets. Each subset independently scores samples in this subset. After averaging results across 10 random seeds, Table VIII shows minimal performance degradation: AC drops by less than $0.4\%$ and AS declines by less than $1.0\%$. This demonstrates consistent effectiveness even with limited data.
IV-D7 Effect of smaller batch size
In Sec. IV-D6, we partition all unlabeled samples into $g\in\{1,2,3\}$ subsets, where mutual scoring is performed only within each subset. To further investigate the impact of reducing the number of samples that simultaneously participate in mutual scoring, we decrease the batch size to less than 10. Table VII reports the results. When the batch size is extremely small (e.g., 5), our MuSc‑V2 shows limited degradation in AC metrics, with a maximum drop of 1.7%. For AS metrics, the performance decreases by 2.4% on the Eyecandies dataset and 4.3% on the MVTec 3D‑AD dataset. This is because, with a batch size of only 5, our interval average operation with a 30% range involves only one sample in scoring, introducing instability. Despite this, the overall degradation remains bounded.
TABLE VII: The performance of a smaller batch size for mutual scoring. We report the AC and AS results. All metrics are in $\%$.
Dataset Batch size F1-max-cls AP-cls F1-max-seg AP-seg
MVTec 3D-AD All 93.0 96.8 54.6 54.7
9 92.1($\downarrow$0.9) 96.0($\downarrow$0.8) 53.3($\downarrow$1.3) 53.0($\downarrow$1.7)
7 91.9($\downarrow$1.1) 95.8($\downarrow$1.0) 51.7($\downarrow$2.9) 51.4($\downarrow$3.3)
5 91.8($\downarrow$1.2) 95.1($\downarrow$1.7) 50.9($\downarrow$3.7) 50.4($\downarrow$4.3)
Eyecandies All 82.9 86.1 44.7 41.8
9 82.8($\downarrow$0.1) 85.9($\downarrow$0.2) 43.9($\downarrow$0.8) 40.5($\downarrow$1.3)
7 82.5($\downarrow$0.4) 85.1($\downarrow$1.0) 43.5($\downarrow$1.2) 40.2($\downarrow$1.6)
5 81.3($\downarrow$1.6) 85.0($\downarrow$1.1) 43.4($\downarrow$1.3) 39.4($\downarrow$2.4)
IV-D8 Sensitivity analysis of hyperparameters
IV-D9 Inference time
Table VIII compares inference times on an NVIDIA RTX 3090 GPU (excluding I/O) across different zero-shot methods. Since the MVTec 3D-AD’s product range is from 100 to 159, we use 150 samples for mutual scoring in MuSc-V2 and MuSc. Our MuSc-V2 outperforms its previous version in both accuracy and speed for 2D tasks. For 3D and multimodal cases, our directly point cloud processing proves significantly faster than PointAD [66], which requires more than 30s per sample for multi-view rendering. While subset partitioning further accelerates MuSc-V2, the 722.6ms feature extraction by Point-MAE [40] remains a bottleneck.
TABLE VIII: Per sample inference time of our MuSc-V2 and other zero-shot methods. We divide the MVTec 3D-AD dataset into $g$ subsets.